OpenStudy (anonymous):

Determine the slope of the line. 9x + 6y − 24 = 0

OpenStudy (displayerror):

Rearrange into point-slope form: \[y = mx + b\] \[6y = -9x + 24\] \[y = -\frac{9}{6} x + 4\] In point-slope form, m represents your slope and b represents your y-intercept. Looking at the last equation, m = -9/6, so that would be your slope.